OpenStudy (anonymous):

If a conditional statement is true, then its contrapositive A . true B. false C. both true and false

OpenStudy (anonymous):

Conditional statement 1: I eat if I am hungry. (I am hungry being the condition) Contrapositive? Yes. I am hungry if I eat. Mathematical: P(A) = (A n B) / P(B) ==> (AnB) = P(A)P(B) P(B) = (A n B) / P(A) ==> (AnB) = P(A)P(B) Undoubtely true.

OpenStudy (anonymous):

thank you